OpenStudy (anonymous):

How could I find a side of a right triangle that gives me the length of the hypotenuse and two of the angles, but doesn't give me sides a or b? (will give medal & fan)

OpenStudy (anonymous):

Specifically, the right triangle i'm working with has a hypotenuse of 25, and two of the angles are 90 degrees (obviously) and 55 degrees... which would make the other angle 35 degrees.

OpenStudy (anonymous):

But I need to find side b of the triangle.

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (whpalmer4):

Do you know about the trigonometric ratios?

OpenStudy (anonymous):

Not exactly... I might, but I need a refresher.

OpenStudy (whpalmer4):

SOH CAH TOA Sin of an angle = Opposite / Hypotenuse Cos of an angle = Adjacent / Hypotenuse Tan of an angle = Opposite / Adjacent We know the angles, and we know the hypotenuse

OpenStudy (whpalmer4):

So we can say, for example, that \[\sin 35 ^\circ = \frac{a}{25}\]

OpenStudy (anonymous):

Oh okay, I see what you did there.

OpenStudy (whpalmer4):

Then \[25\sin 35^\circ = a\]

OpenStudy (anonymous):

so a = 14.34 (rounded to two decimal places)?

OpenStudy (whpalmer4):

Indeed!

OpenStudy (whpalmer4):

then you could find the other one with Pythagorean theorem or one of the other trig ratios

OpenStudy (anonymous):

Okay so when all is said and done b = 2.04 (rounded)?

OpenStudy (anonymous):

*3.04

OpenStudy (anonymous):

b + 14.34^2 = 25^2 b + 205.6356 = 625 b = 625/205.6356 b = 3.04

OpenStudy (whpalmer4):

No, that's not quite right... You are using a little-known variation of the Pythagorean theorem there: \[a^2+b = c^2\]:-)

OpenStudy (whpalmer4):

As a reasonableness check, look at your triangle: 55 and 35 degrees aren't quite equal, but they are both pretty close to 45 degrees, which would give you an isosceles triangle with a = b....you've got something more like a = 5b

OpenStudy (anonymous):

so wait, a = 14.34 but wouldn't b be larger than that?

OpenStudy (anonymous):

I'm confused now lol

OpenStudy (whpalmer4):

yes, \(b\) is larger than \(a\). If we had a 45/45/90 triangle, a would equal b. I'm saying the angles are close enough to equal that the sides should be much closer to each other in length. A triangle with the sides you computed is impossible — think about it! The other two sides added together must be at least the length of the hypotenuse, right? :-)

OpenStudy (anonymous):

So then... a^2 + b^2 = c^2 14.34^2 + b^2 = 25^2 Am I doing this right so far?

OpenStudy (whpalmer4):

yes!

OpenStudy (whpalmer4):

before you did\[14.34^2 + b = 25^2\]

OpenStudy (anonymous):

okay so 205.6356 + b^2 = 625 b^2 = 625/205.6356 b^2 = 3.04 b = 1.74 ... still not working the way i want it to...

OpenStudy (anonymous):

ohhh wait i get it i'm dividing instead of subtracting

OpenStudy (anonymous):

b = 20.48

OpenStudy (whpalmer4):

Yes, that's the right answer!

OpenStudy (anonymous):

yay! thank you!